# PixelWorld: Towards Perceiving Everything as Pixels

### v2

@@ Abstract
Recent agentic language models increasingly need to interact directly with real-world environments containing intertwined visual and textual information through raw camera pixels, rather than relying on separate image and tokenized text processing, underscoring the necessity of a unified perception paradigm. To close this gap, we explore this idea through Perceive Everything as Pixels (PEAP) and release PixelWorld, a benchmark that renders natural-language, tabular, mathematical and diagrammatic inputs into a single pixel space. Experiments show that PEAP attains competitive accuracy on semantic-understanding tasks, indicating that a vision transformer can capture global textual semantics without explicit tokens. In contrast, reasoning-intensive benchmarks (math and code) exhibit sharp performance drops; however, Chain-of-Thought prompting partially mitigates this gap, hinting that explicit reasoning traces compensate for the missing token structure. We also find that when visual and textual information are closely integrated, representing everything as pixels reduces preprocessing complexity and avoids misalignment issues that often arise in separate pipelines. PixelWorld therefore serves as a practical benchmark for evaluating unified vision–language models and supports broader exploration of PEAP across diverse tasks.
PixelWorld: Towards Perceiving Everything as Pixels
Zhiheng Lyu1,2 Xueguang Ma1 Wenhu Chen1,2 1University of Waterloo  2Vector Institute, Toronto {z63lyu,x93ma,wenhuchen}@uwaterloo.ca
@@ 1 Introduction
Figure 1: PEAP framework: we investigate the possibility of perceive everything as pixels. This framework aligns better with human perception reducing the need for excessive pre-processing. Evaluated on our benchmark PixelWorld, PEAP boosts performance on multimodal tasks (e.g., websites, slides, documents) but struggles with complex, text-centric tasks (e.g., reasoning and coding). Larger models achieve better transferability between pixel- and token-based performance compared to smaller ones. We also observed that text and images exhibit similar attention patterns, and reduced the overhead of model reasoning through patch pruning by PEAP-Fast.
@@
In recent years, large vision-language models (L-VLMs) (Wang et al., 2024a; OpenAI, 2025; Team, 2024) have achieved impressive performance across a wide range of real-world tasks. These models typically process visual inputs as pixels and textual inputs as discrete tokens—two distinct modalities that are handled separately. However, such modality-specific processing leads to a fragmented understanding of multimodal inputs and increases the complexity of engineering pipelines. This separation becomes particularly problematic in modern agent-based systems such as Computer Agents (Zheng et al., 2024; Koh et al., 2024) and Embodied Agents (Tellex et al., 2020; Driess et al., 2023), which are increasingly expected to perform complex real-world tasks including navigation in physical environments (Elnoor et al., 2024), booking flights (Chen et al., 2024a), and repairing software bugs on platforms like GitHub (Yang et al., 2024). These tasks involve deeply intertwined visual and textual information, where decoupled tokenization and perception modules can result in high preprocessing overhead (Xie et al., 2024; Koh et al., 2024) and degraded performance due to information loss and layout inconsistencies (Dagan et al., 2024; Chai et al., 2024).
@@
Due to these limitations, we propose a unified perception paradigm: Perceive Everything as Pixels (PEAP). In this paradigm, both text and visual inputs are treated uniformly in the pixel space, allowing a vision-language model (VLM) to jointly model multimodal inputs without separate tokenization or modality-specific encoders. To identify the benefits and challenges of this paradigm, we introduce PixelWorld, a comprehensive benchmark suite designed to assess how well VLMs perform on existing benchmarks under the PEAP setting.
@@
In PixelWorld, we select 10 representative commonly used benchmarks, covering a diverse range of modalities and task scenarios. For each dataset, we construct both traditional token-based and pixel-based (PEAP) input formats using image synthesis and OCR techniques (see Table 1). We then evaluate vision–language models of varying scales, from Qwen2VL-2B to GPT-4o. Cross-modal evaluation in Section 2 yields three overarching insights: Insight 1: in intrinsically multimodal settings such as website rendering, slide comprehension, and document understanding, PEAP eliminates OCR noise and consistently boosts performance; Insight 2: pixelising inputs for reasoning intensive tasks such as math and code incurs marked accuracy drops, yet the gap narrows as model capacity grows, suggesting scale is critical for cross-modal transfer; Insight 3: larger models demonstrate superior instruction-following and long-context reasoning across modalities, while smaller models struggle, emphasizing the need for scale-aware training under the pixel paradigm.
@@
To further understand these findings, we conduct additional analyses from three perspectives: (1) Representation analysis: We visualize the attention patterns of Qwen2VL-7B and find consistent global structures between token- and pixel-based inputs, suggesting that vision encoders can serve as universal tokenizers. (2) Efficiency optimization: We measure inference latency and show that while PEAP increases processing time due to input size, our proposed PEAP-Fast algorithm prunes blank patches and achieves up to 80% speedup without accuracy loss. (3) Prompt sensitivity: We explore input prompting strategies and find that Chain-of-Thought (CoT) boosts performance on PEAP more effectively than standard input.
@@
    PixelWorld: We introduce a unified benchmark that converts text, structural, and multimodal datasets into pixels, providing a direct stress test and diagnostic tool for PEAP. The dataset and evaluation code are publicly released to foster a more holistic yardstick for future L-VLMs research and to ease multimodal data collection.
@@
    Task–scale insights: PEAP consistently improves layout-heavy or intrinsically multimodal tasks (e.g., website and document understanding) but degrades on reasoning- and code-centric benchmarks; the performance gap diminishes as model size increases, highlighting scale as a key factor for transferability.
@@
    Efficiency & interpretability: We propose PEAP-Fast, which removes blank pixel patches to achieve up to a $3\times$3 × latency reduction without harming accuracy. Attention visualizations reveal similar global patterns between pixel- and token-based models, suggesting that vision encoders can act as a universal multimodal tokenizer.
@@
GLUE Wang (2018) 59,879 Natural language understanding Synthesis test
SuperGLUE Sarlin et al. (2020) 19,294 Natural language understanding Synthesis test
MMLU-Pro Wang et al. (2024b) 12,032 Domain knowledge and reasoning Synthesis test
ARC Clark et al. (2018) 3,548 Science question answering Synthesis test
GSM8K Cobbe et al. (2021) 1,319 Math problem solving Synthesis test
MBPP Austin et al. (2021) 757 Programming tasks Synthesis test
@@
TableBench Wu et al. (2024) 888 Table data understanding and analysis Synthesis test
@@
MathVerse Zhang et al. (2025) 788 Math and visual reasoning Natural test
MMMU-Pro Yue et al. (2024) 1,730 Multimodal reasoning Synthesis test
@@ 2 Datasets
2.0.0.0.1 Image Data Synthesis
@@ 3 Experiments
Metrics For question-answering tasks such as WikiSS-QA, SlidesVQA, and TableBench, we adopt ROUGE-L as our primary metric, as it effectively captures the alignment between generated answers and ground truth by measuring the longest common subsequence. For classification benchmarks, including MMLU-Pro, GLUE, SuperGLUE, ARC, and MathVerse, we use accuracy, which directly reflects the model’s performance in selecting correct options. For GLUE and SuperGLUE, we follow their standard evaluation protocols, utilizing task-specific metrics such as Matthews correlation, F1 score, and Pearson correlation. For the code generation task MBPP, we evaluate performance using the pass@1 rate, which measures whether the generated code successfully passes all test cases. For the mathematical reasoning dataset GSM8K, we employ exact match accuracy, as these problems require precise numerical answers. For the visualization subtask of TableBench, following the original codebase, we treat it as a code generation task and evaluate the correctness of the generated visualizations.
@@
Model Selection To validate PixelWorld, we selected a diverse set of vision-language models (VLMs) with varying scales to ensure the robustness and generalizability of our findings. It also allowed us to analyze the behavior of models across different sizes. We evaluated several widely used vision-language models (VLMs), including Qwen2VL-2B Wang et al. (2024a), Phi-3.5-3.2B Abdin et al. (2024), Qwen2VL-7B Wang et al. (2024a), Gemini-Flash Team (2024), and GPT-4o OpenAI (2025).
@@ 3.3 Multimodal Input
Figure 6: Last Layer Attention Heatmap on Qwen2VL-7B between token-based (left) and pixel-based (right) inference.
@@ 4.1 Q1: Does PEAP have the same attention?
To investigate whether VLMs behaves similarly on textual and image inputs, we visualized the Average Attention of Qwen2-VL-7B’s final layer using a heatmap (see Figure 6). Concretely, we examined its responses on a SuperGLUE BoolQ example, comparing the model’s attention maps for text-based versus image-based inference.
As shown in Figure 6, the model largely focuses on task-relevant elements such as the question prompt (“will there be a sequel …”), the key words in the passage (e.g., “film”, “starring”, “Alice”), and the required answer format (“Answer: True/False”). This holds true across both textual and visual representations, indicating Qwen2-VL-7B exhibits comparable attention patterns irrespective of input modality. However, we also observe that certain blank patches in the image-based input can receive disproportionately high attention. This suggests that while the visual encoder parallels the text encoder in many respects, it still has redundancy.
@@ 4.3 Q3: Is PEAP sensitive to the prompting method?
Massive experimental results in Section 2 show that the performance gap between image and text inputs still exists, potentially due to domain gaps in datasets or insufficient instruction following in image inputs. To address this, we applied CoT-style prompts to the SuperGLUE dataset to enhance cross-domain instruction following (Table 4). Notably, Qwen2VL-7B showed significant improvements in tasks where image input underperformed compared to text input, such as CB and RTE. Overall, CoT prompts improved image input performance by 2.58%, surpassing the 0.3% improvement observed for text input.
@@ 5 Related Work
5.0.0.0.1 Multimodal Large Language Models and Benchmarks
Recent progress in multimodal AI has led to the development of models like GPT-4o OpenAI (2025), Gemini Team (2024), and Claude-3.5 Anthropic (2025), which integrate vision-based training to improve instruction-following capabilities. Benchmarks for these models have evolved from task-specific datasets, such as VQA Agrawal et al. (2016) and DocVQA Mathew et al. (2021), to more comprehensive evaluations, including MMMU-Pro Yue et al. (2024), MMBench Liu et al. (2024), and MegaBench Chen et al. (2024b). However, most current research focuses on the semantic understanding of visual content, with only a few benchmarks—such as MathVerse Zhang et al. (2025) and MMMU-Pro Yue et al. (2024)—addressing text recognition and comprehension within images. Our work shifts the focus towards evaluating how well large language models understand language through visual input compared to traditional token-based input.
5.0.0.0.2 Screenshot LMs
Recent studies have demonstrated that pretraining on synthetic screenshots can enable vision-language models (VLMs) to achieve performance comparable to that of BERT on language modeling tasks Lee et al. (2022); Rust et al. (2023); Gao et al. (2024). This approach allows models to better capture text structures without relying on OCR-based methods. Furthermore, our analysis highlights a performance gap between existing VLMs on vision-based tasks and their text-only counterparts, particularly in the absence of relevant pretraining. Interestingly, in certain scenarios, VLMs perform as well as or even better than text-only models, underscoring the potential of this research direction. In the context of document retrieval, recent advancements Faysse et al. (2024); Ma et al. (2024) have shown that large-scale pretraining on screenshots can outperform traditional OCR-based methods, further reinforcing the advantages of vision-language pretraining.
5.0.0.0.3 Language Tokenization
Tokenization methods, such as Byte Pair Encoding (BPE) Shibata et al. (1999); Sennrich et al. (2016), are widely used in language modeling, but recent studies suggest that they may not always be optimal. For instance, MegaByte Yu et al. (2023) demonstrated that fixed-length tokenization can improve both computational efficiency and cross-modal capabilities. Similarly, BLT Pagnoni et al. (2024) proposed entropy-based tokenization, while LCM team et al. (2024) emphasized the benefits of processing higher-level semantic concepts rather than individual tokens. Inspired by these approaches, we explore whether adaptive image patches can effectively infer textual meaning. At a higher level, we investigate the unification of text and image inputs into a shared representation space, enabling reasoning through abstract semantic concepts rather than traditional token-based methods.

### v3

# Zhiheng Lyu1,2 Xueguang Ma1 Wenhu Chen1,2

@@ Abstract
Recent agentic language models increasingly need to interact with real-world environments that contain tightly intertwined visual and textual information, often through raw camera pixels rather than separately processed images and tokenized text. This shift highlights the need for a unified perception paradigm. To investigate this idea, we explore Perceive Everything as Pixels (PEAP) and introduce PixelWorld, a benchmark that renders natural-language, tabular, mathematical, and diagrammatic inputs into a shared pixel space. Experiments across multiple benchmarks show that PEAP achieves comparable performance to token-based approaches on semantic understanding tasks, suggesting that vision transformers can partially capture global textual semantics without explicit tokenization. In contrast, reasoning-intensive tasks such as mathematics and code show notable performance degradation, although Chain-of-Thought prompting helps mitigate this gap by compensating for missing symbolic structure. We further find that when visual and textual information are closely integrated, representing everything as pixels simplifies preprocessing and avoids cross-modal misalignment. PixelWorld thus provides a systematic and practical framework for evaluating unified vision–language models and facilitates further exploration of pixel-based multimodal learning.
@@ 1 Introduction
In recent years, large vision–language models (L-VLMs) (Wang et al., 2024a; OpenAI, 2025; Gemini, 2024) have achieved remarkable progress across diverse real-world tasks. However, these models still rely on distinct processing pipelines for different modalities—treating images as pixels and text as discrete tokens. Such disjoint tokenization introduces a representation mismatch between vision and language, which hampers unified multimodal understanding and complicates system design. As recent works (Zheng et al., 2024; Koh et al., 2024; Tellex et al., 2020; Driess et al., 2023) push toward agentic systems capable of perceiving and acting in complex environments—from physical navigation (Elnoor et al., 2024) and travel booking (Chen et al., 2024a) to code repair on GitHub (Yang et al., 2024)—this mismatch becomes increasingly consequential. In such intertwined visual–textual settings, maintaining separate tokenization and perception modules not only incurs high preprocessing overhead (Xie et al., 2024; Koh et al., 2024) but also leads to information loss and layout inconsistencies (Dagan et al., 2024; Chai et al., 2024), ultimately limiting the scalability and robustness of multimodal agents.
@@
To address these limitations, we explore a unified perception approach: Perceive Everything as Pixels (PEAP). Building on earlier efforts that considered pixel-based representations (Singh et al., 2024; Zhang et al., 2024), we systematically examine how representing both text and visual inputs uniformly in pixel space affects model behavior. In this paradigm, a vision–language model (VLM) jointly models multimodal inputs without requiring separate tokenization or modality-specific encoders. To better understand the benefits and challenges of this approach, we introduce PixelWorld, a comprehensive benchmark suite designed to evaluate how well existing VLMs perform under the PEAP setting.
Figure 1: Overview of the PEAP Framework. PEAP (Perceive Everything as Pixels) unifies text, structural, and visual inputs into a single pixel space, where a Vision Transformer (ViT) encodes the pixels and a language decoder performs reasoning. Both components are enclosed within the dashed box to indicate that they jointly form a vision–language model (VLM). By eliminating modality-specific preprocessing such as OCR and tokenization, PEAP better aligns with human perception and reduces cross-modal engineering overhead.
Figure 2: Key Findings on the PixelWorld Benchmark. Evaluated across text-only, structural, and multimodal settings (§2, §3), PEAP shows four major insights: (1) Modality Trend: consistent gains on layout-heavy and multimodal tasks such as websites, slides, and documents; (2) Task Complexity: performance degradation on reasoning- and code-centric benchmarks (see §3.1–§3.2); (3) Transferability by Scale: larger VLMs (e.g., GPT-4o, Gemini-Flash) exhibit smaller pixel–token gaps; and (4) Attention and Efficiency: text and image inputs show similar global attention patterns, while the proposed PEAP-Fast reduces up to 80% of computation overhead (§4.2).
@@
In PixelWorld, we select 10 widely used benchmarks covering a diverse range of modalities and task scenarios. For each dataset, we construct both traditional token-based and pixel-based (PEAP) input formats using image synthesis and OCR techniques (see Table 1). We then evaluate vision–language models of varying scales, from Qwen2VL-2B to GPT-4o. Cross-modal evaluation in Section 3 reveals three overarching insights: Insight 1: In intrinsically multimodal settings such as website rendering, slide comprehension, and document understanding, PEAP consistently mitigates OCR noise and yields stronger performance; Insight 2: For reasoning-intensive tasks such as mathematics and code, pixelization leads to noticeable accuracy drops, though the gap narrows as model capacity increases—suggesting that scale plays a key role in enabling cross-modal transfer; and Insight 3: Larger models exhibit more robust instruction-following and long-context reasoning across modalities, whereas smaller models struggle, highlighting the importance of scale-aware training under the pixel-based paradigm.
@@
To further interpret these findings, we conduct three complementary analyses. (1) Representation analysis: We visualize the attention patterns of Qwen2VL-7B and observe broadly similar global structures between token- and pixel-based inputs, suggesting that certain aspects of language modeling behavior may transfer into the visual space, though not implying full equivalence. (2) Efficiency optimization: We measure inference latency and find that while PEAP increases computational cost due to larger input resolution, our proposed PEAP-Fast algorithm effectively prunes blank patches, achieving up to 80% speedup with negligible accuracy degradation. (3) Prompt sensitivity: We study prompting strategies and find that Chain-of-Thought (CoT) reasoning yields more consistent gains under the pixel-based representation compared to standard direct prompting, indicating potential synergies between reasoning supervision and visual encoding.
@@
    PixelWorld Benchmark: We present a unified benchmark that transforms text, structural, and multimodal datasets into pixel space, offering a direct and reproducible framework to evaluate the trade-offs between pixel- and token-based modeling. The benchmark and code are publicly released to facilitate standardized comparison and future research on multimodal perception.
@@
    Task–scale insights: Through large-scale evaluation, we show that PEAP improves layout-heavy or intrinsically multimodal tasks (e.g., website and document understanding) while reducing accuracy on reasoning- or code-centric tasks. The performance gap consistently narrows with model scale, underscoring the role of capacity in enabling cross-modal transfer.
@@
    Efficiency & interpretability: We propose PEAP-Fast, an inference-time pruning strategy that removes blank pixel patches, achieving up to a $3\times$ latency reduction with minimal loss in accuracy. Attention visualizations reveal partially shared global structures across modalities, providing an interpretable perspective on how visual encoders approximate token-level reasoning behavior.
@@
GLUE (Wang, 2018) 59,879 Natural language understanding Synthesis test
SuperGLUE (Sarlin et al., 2020) 19,294 Natural language understanding Synthesis test
MMLU-Pro (Wang et al., 2024b) 12,032 Domain knowledge and reasoning Synthesis test
ARC (Clark et al., 2018) 3,548 Science question answering Synthesis test
GSM8K (Cobbe et al., 2021) 1,319 Math problem solving Synthesis test
MBPP (Austin et al., 2021) 757 Programming tasks Synthesis test
@@
TableBench (Wu et al., 2024) 888 Table data understanding and analysis Synthesis test
@@
MathVerse (Zhang et al., 2025) 788 Math and visual reasoning Natural test
MMMU-Pro (Yue et al., 2024) 1,730 Multimodal reasoning Synthesis test
@@ 2 Datasets
Image Data Synthesis
@@ 3 Experiments
Metrics For QA tasks (WikiSS-QA, SlidesVQA, TableBench), we use ROUGE-L, which measures the longest common subsequence between prediction and reference to approximate answer overlap. We choose it for convenience and comparability, and expect other semantic metrics (e.g., BERTScore, LLM judges) to show similar trends. For classification benchmarks, including MMLU-Pro, GLUE, SuperGLUE, ARC, and MathVerse, we use accuracy, which directly reflects the model’s performance in selecting correct options. For GLUE and SuperGLUE, we follow their standard evaluation protocols, utilizing task-specific metrics such as Matthews correlation, F1 score, and Pearson correlation. For the code generation task MBPP, we evaluate performance using the pass@1 rate, which measures whether the generated code successfully passes all test cases. For the mathematical reasoning dataset GSM8K, we employ exact match accuracy, as these problems require precise numerical answers. For the visualization subtask of TableBench, following the original codebase, we treat it as a code generation task and evaluate the correctness of the generated visualizations.
@@
Model Selection To validate PixelWorld, we selected a diverse set of vision-language models (VLMs) with varying scales to ensure the robustness and generalizability of our findings. It also allowed us to analyze the behavior of models across different sizes. We evaluated several widely used vision-language models (VLMs), including Qwen2VL-2B (Wang et al., 2024a), Phi-3.5-3.2B (Abdin et al., 2024), Qwen2VL-7B (Wang et al., 2024a), Gemini-Flash (Gemini, 2024) [^1] We use gemini-1.5-flash-002, which was the latest available version during this study., and GPT-4o (OpenAI, 2025).
@@ 3.3 Multimodal Input
Figure 7: Last Layer Attention Heatmap on Qwen2VL-7B between token-based (left) and pixel-based (right) inference. Although the overall attention intensity on image inputs is generally lower, both modalities exhibit highly similar attention patterns.
@@ 4.1 Q1: Does PEAP have the same attention?
To examine whether VLMs attend to similar regions when processing textual and image inputs, we visualize the average attention map of the final layer in Qwen2-VL-7B using a heatmap (Figure 7). Specifically, we analyze the model’s behavior on a BoolQ example from SuperGLUE, comparing its attention patterns under text-based and image-based inference. Similar attention behaviors are observed across different datasets; more examples are shown in Appendix C.
Computation of Heatmaps. The attention heatmaps are computed during greedy decoding by averaging the last-layer attention across tokens. For multi-head attention, we apply a simple mean across heads. Formally, the attention weight for position $i$ is given by:
$\text{Heatmap}(i)\;=\;\frac{1}{H}\sum_{h=1}^{H}\big|A^{(L)}_{h}[t,i]\big|,\quad i\in[s,e)$
where $H$ denotes the number of attention heads, $L$ is the index of the last layer, $A^{(L)}_{h}[t,i]$ represents the attention weight from token $t$ to token $i$ in head $h$, and $[s,e)$ corresponds to the decoding range during greedy decoding.
Observations. As shown in Figure 7, Qwen2-VL-7B consistently focuses on task-relevant elements such as the question prompt (“will there be a sequel …”), salient passage keywords (e.g., “film”, “starring”, “Alice”), and the expected answer format (“Answer: True/False”). This pattern remains stable across both textual and visual representations, suggesting that the model exhibits largely comparable attention behavior regardless of input modality. However, we also observe that certain blank patches in the image-based inputs occasionally receive disproportionately high attention weights, indicating that while the visual encoder aligns closely with the text encoder, it still introduces redundant activations.
@@ 4.3 Q3: Is PEAP sensitive to the prompting method?
In Section 3, results on purely textual synthesis tasks show that image-based inputs consistently underperform text inputs, likely due to dataset domain gaps and weaker instruction following on image representations. To address this, we applied CoT-style prompts to the SuperGLUE dataset to enhance cross-domain instruction following (Table 4). Notably, Qwen2VL-7B showed significant improvements in tasks where image input underperformed compared to text input, such as CB and RTE. Overall, CoT prompts improved image input performance by 2.58%, surpassing the 0.3% improvement observed for text input.
@@ 5 Related Work
Multimodal Large Language Models and Benchmarks
Recent progress in multimodal AI has led to the development of models like GPT-4o (OpenAI, 2025), Gemini (Gemini, 2024), and Claude-3.5 (Anthropic, 2025), which integrate vision-based training to improve instruction-following capabilities. Benchmarks for these models have evolved from task-specific datasets, such as VQA (Agrawal et al., 2016) and DocVQA (Mathew et al., 2021), to more comprehensive evaluations, including MMMU-Pro (Yue et al., 2024), MMBench (Liu et al., 2024), and MegaBench (Chen et al., 2024b). However, most current research focuses on the semantic understanding of visual content, with only a few benchmarks—such as MathVerse (Zhang et al., 2025) and MMMU-Pro (Yue et al., 2024)—addressing text recognition and comprehension within images. Our work shifts the focus towards evaluating how well large language models understand language through visual input compared to traditional token-based input.
Screenshot LMs
Recent studies have demonstrated that pretraining on synthetic screenshots can enable vision-language models (VLMs) to achieve performance comparable to that of BERT on language modeling tasks (Lee et al., 2022; Rust et al., 2023; Gao et al., 2024). This approach allows models to better capture text structures without relying on OCR-based methods. Furthermore, our analysis highlights a performance gap between existing VLMs on vision-based tasks and their text-only counterparts, particularly in the absence of relevant pretraining. Interestingly, in certain scenarios, VLMs perform as well as or even better than text-only models, underscoring the potential of this research direction. In the context of document retrieval, recent advancements (Faysse et al., 2024; Ma et al., 2024) have shown that large-scale pretraining on screenshots can outperform traditional OCR-based methods, further reinforcing the advantages of vision-language pretraining.
Language Tokenization
Tokenization methods, such as Byte Pair Encoding (BPE) (Shibata et al., 1999; Sennrich et al., 2016), are widely used in language modeling, but recent studies suggest that they may not always be optimal. For instance, MegaByte (Yu et al., 2023) demonstrated that fixed-length tokenization can improve both computational efficiency and cross-modal capabilities. Similarly, BLT (Pagnoni et al., 2024) proposed entropy-based tokenization, while LCM (team et al., 2024) emphasized the benefits of processing higher-level semantic concepts rather than individual tokens. Inspired by these approaches, we explore whether adaptive image patches can effectively infer textual meaning. At a higher level, we investigate the unification of text and image inputs into a shared representation space, enabling reasoning through abstract semantic concepts rather than traditional token-based methods.